OpenStudy (anonymous):

write an equation of the line containing the given point and parrallel to the given line. xpress in y=mx+b. (-3,3) 2x= 3y+8. HELP ME

OpenStudy (anonymous):

parallel lines have same slope

OpenStudy (anonymous):

use point slope to find the parallel equation and then convert to slope-intercept point slope: (y-y1)=m(x-x1)

OpenStudy (anonymous):

how do you get it to slope- intervept form?

OpenStudy (anonymous):

y = mx + b

OpenStudy (anonymous):

which would be m?

OpenStudy (anonymous):

2x= 3y+8 3y = 2x - 8 y = (2/3) x - (8/3) so slope (m) = 2/3

OpenStudy (anonymous):

is this the equation. y-3= 2/3 (x-(-3)) for point slope ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

simplify it if you wish

OpenStudy (anonymous):

how do i simplify ?

OpenStudy (anonymous):

Thanks